OpenStudy (anonymous):

Simplify: If P(x) = 3x2 - 7x + 4 and Q(x) = -5 + 8x

OpenStudy (anonymous):

P(X)=(X-1) (3x-4) 3x^2+4=p(x)+7x

OpenStudy (anonymous):

this is what the question is and i dont really understand it :/ If P(x) = 3x2 - 7x + 4 and Q(x) = -5 + 8x a. P(-6) b. P(x) + Q(x) c. P(x) – Q(x)